# Tensor Network Computations That Capture Strict Variationality, Volume Law Behavior, and the Efficient Representation of Neural Network States

### v2

@@ Introduction.
Tensor networks represent the amplitude of a quantum state on an $L$ site lattice, denoted $\langle\vec{n}|\Psi\rangle\equiv\Psi(\vec{n})\equiv\Psi(n_{1},n_{2},\ldots,n_{L})$, as the contraction of a set of tensors. As a simple example, we consider $L$ tensors connected in a graph (Fig. 1a), each associated with a lattice site Hilbert space; this is known as a projected entangled pair state (PEPS) [27]. Each tensor contains $dD^{m}$ elements ($m$ is the number of bonds which connect tensors to adjacent tensors): $d$ is the local Hilbert space dimension and $D$ controls the expressivity. For tree graphs, the resulting contraction over the shared tensor bonds can be computed exactly in $O(L)\mathrm{poly}(D)$ time [28]. For general graphs, however, both $\langle\vec{n}|\Psi\rangle$ and $\langle\Psi|\hat{O}|\Psi\rangle$ cannot be exactly contracted efficiently: the cost is exponential in the tree-width of the graph [21]. Approximate contraction algorithms have thus been introduced whereby during the contraction process, the large intermediate tensors that are formed are compressed, via singular value decomposition (SVD), to a controlled size, $O(\mathrm{poly}(\chi))$, where $\chi$ is termed the auxiliary bond dimension [2, 29]. Only in the limit of $\chi\to\infty$ is the contraction exact. The resulting approximate contraction has a cost $O(L)\mathrm{poly}(D)\mathrm{poly}(\chi)$, with an error with respect to the exact contraction controlled by $\chi$.
@@
From the perspective of approximate contraction, tensor network computations should seek to be converged with respect to $\chi$, otherwise the results are not meaningful. For example, for a Hamiltonian $\hat{H}$, if $\langle\Psi|\hat{H}|\Psi\rangle/\langle\Psi|\Psi\rangle$ is computed by approximate contraction, with insufficient $\chi$ the energy may violate the variational theorem. The requirement that computations should use a sufficiently large $\chi$ (such that the evaluation of $\langle\Psi|\hat{O}|\Psi\rangle$ or $\langle\vec{n}|\Psi\rangle$ is no longer changing significantly with $\chi$) places restrictions on the types of graphs that can be used to define the tensor network. For example, in many-body physics simulations, it is believed that only highly structured graphs with limited entanglement, such as (near-)area law states, support an efficient contraction scheme. Because of the latter requirement, it has been argued that, in practical use, tensor networks cannot replicate the expressivity of other flexible types of many-body ansatz, such as neural network quantum states [30, 31, 32, 33, 34, 35, 36].
@@
In the current work, we observe that the above limitations on tensor network computations arise due to an assumption, namely that the tensor network ansatz is only meaningfully defined in the exact contraction limit. We can, however, adopt an alternative perspective, namely, to consider the approximate contraction scheme for any $\chi$ as itself part of the tensor network ansatz for the amplitude $\Psi(\vec{n})$. This view is natural and in some ways obvious when using tensor networks in variational Monte Carlo (VMC) calculations [37, 38, 39, 40, 41], but we will argue that this is not a manifestation of tensor networks in a particular numerical algorithm, but rather a change of perspective that alters and broadens the definition and utility of tensor networks. We refer to the larger set of parameterized many-body states that arise from “consistent” contraction schemes (to be defined later) as tensor network functions. This wider class of states removes many restrictions that are assumed to apply to tensor networks. We demonstrate that tensor network functions provide strictly variational estimates of the energy, allow for practical tensor network computations on graphs that are highly challenging to contract accurately, including those that describe volume law physics, and expand the set of states that can be efficiently represented by tensor networks to cover other classes of ansatz, such as neural network quantum states.
@@
Figure 1: (a) PEPS on a general graph with $L$ site tensors. (b) PEPS on a $4\times 4$ square lattice. (c) Insertion of a pair of isometries for approximate tensor contractions; the isometries can be obtained with standard tensor network techniques by SVD [2, 3]. (d) Approximate PEPS contractions by inserting isometries for a given configuration $|\vec{n}\rangle$. (e) For a given input vector $|\vec{n}\rangle=(n_{1},n_{2},...,n_{L})$, a tensor network function outputs a unique value $\Psi(\vec{n})$.
@@
A tensor network function (TNF) is a consistent contraction of the tensor network representation of the amplitude $\Psi(\vec{n})$. To illustrate the basic idea, we first consider a standard PEPS for a two-dimensional lattice, illustrated in Fig. 1b. Because the graph has loops, the PEPS amplitude cannot be computed efficiently according to Fig. 1b. Instead, in practice, the amplitude is defined by a different tensor network illustrated in Fig. 1d. In this figure, isometries (triangles) with bond dimension $\chi$ have been inserted, where the values of the isometries are determined by an SVD of the contraction of neighbouring tensors (see Fig. 1c) [2, 3]. The final amplitude is obtained by contracting the tensors and isometries together. The pattern of isometries used here corresponds to a specific approximate contraction method (boundary contraction [27, 2]) and other choices can be made [29, 42], but the essential purpose is to make the contraction of the modified graph containing the original tensors and the isometries efficient.
Note, however, that although the structure of the tensor network with isometries in Fig. 1d now appears pseudo-one-dimensional, this does not mean that its entanglement structure (for $\chi>1$) is reduced to that of a matrix product state (MPS). This is because the isometries are themselves functions of the configuration $|\vec{n}\rangle$, i.e. different values of the isometries are obtained, via SVD, for different configurations. However, we choose to keep the same positions of the isometries in the graph for any configuration (even though the entries are different); we refer to this as using fixed (position) isometries, which defines a consistent contraction. Because of the configuration dependence of the isometry entries, the graph with isometries no longer represents a multi-linear ansatz like the original tensor network state; it is only a multi-linear ansatz in the limit that $\chi\to\infty$ (and the isometries become identities). Instead, it defines a tensor network function: for any given $(D,\chi)$, any $|\vec{n}\rangle$ is mapped to a single value, the amplitude, through an efficient contraction. The efficient TNF representation of $\Psi(\vec{n})$ does not guarantee deterministic efficiency of computing $\langle\Psi|\hat{O}|\Psi\rangle$; but such expectations may be evaluated stochastically by sampling with efficient sample complexity.
@@
VMC algorithms also compute $\Psi(\vec{n})$ via approximate contraction (for a brief review of tensor network VMC (TN-VMC) algorithms, see Refs. [37, 38, 39, 40, 41]) and it is instructive to consider the differences between the standard TN-VMC approximation of $\Psi(\vec{n})$ and the definition of a TNF above. There are two seemingly small, but significant, differences (i) in TN-VMC, $\chi$ is viewed as a parameter (typically $\chi=\mathrm{const}\times D$, with $\mathrm{const}>1$) where $\chi$ is increased to convergence for a fixed $D$, and (ii) the positions of the isometries are not fixed but dependent on the amplitude that is being computed (we refer to this as dynamic isometries), in order to allow for the reuse of computations when computing amplitudes of configurations which differ by only a few substitutions [41].
The most important distinction from TNF is that the use of dynamic isometries in TN-VMC does not lead to a consistent function. In other words, given the same configuration $|\vec{n}\rangle$, $\Psi(\vec{n})$ needs not be the same, as it depends on the prior configuration and which components of the tensor network contraction are being reused. Only in the limit of $\chi\to\infty$ are the amplitudes consistent and the function defined, which underlies the usual need to converge $\chi$ in TN-VMC calculations. In contrast, it is clear that with using fixed isometries, the corresponding TNF is consistent for any $D$ and any $\chi$.
@@
In Fig. 2(a) we show the energy of an optimized PEPS ($D=8$, various $\chi$) TNF for the ground-state of the open-boundary condition (OBC) spin-1/2 square-lattice $32\times 32$ Heisenberg model, compared to that of a PEPS in standard TN-VMC using dynamic isometries, where $\chi$ controls the error of approximate contraction [41]. We see immediately that the PEPS TNF energy is variational for all $\chi$. In contrast, viewing the PEPS amplitude as an (inconsistent) approximate contraction as in standard VMC leads to a non-variational energy, where the non-variationality is typically regarded as the “contraction error”. This behaviour is even more clear for a lattice that is usually considered more difficult to contract. Using the square-lattice $J_{1}-J_{2}$ model with periodic boundary conditions (PBC) as an example, in Fig. 2(b) we show the same comparison between TNF and standard TN-VMC. In addition to the variational versus non-variational nature of the energies, we also see a large variance of the TN-VMC energy arising from the inconsistent definition of amplitudes. This leads to substantial difficulties in the stochastic optimization of the ansatz. Overall, we see that lack of variationality is not intrinsic to approximate contraction: it results from the perspective on the definition of the ansatz.
@@
In Figs. 3(a) and (b), we show results from approximating the ground-state of a spin-$1/2$ square-lattice $16\times 16$ $J_{1}-J_{2}$ Heisenberg model for $J_{2}/J_{1}=0.5$, and an $8\times 8$ Hubbard model with $U/t=8$ and $1/8$ hole doping, using bosonic and fermionic PEPS TNFs [27, 43, 44, 45], respectively. We minimize the energy for many different $(D,\chi)$ combinations using gradient descent, and we plot the computational cost of the amplitude against the obtained accuracy. For a specified relative accuracy, we can then search for the $(D,\chi)$ pair with the lowest computational cost. We see that contrary to the standard situation with tensor network states where one uses $\chi>D$, the most powerful TNFs for a given computational cost have $\chi<D$. For example, for $\sim 0.01$ relative accuracy in the Hubbard model, using $(D,\chi)=(12,6)$ yields the same expressivity as $(D,\chi)=(10,10)$, but with lower cost. In no case is it computationally more efficient to use $\chi>D$.
Figure 3: Time cost (in seconds) versus accuracy of the tensor network function $|\Psi(D,\chi)\rangle$, for the $J_{1}-J_{2}$ model on the $16\times 16$ lattice at $J_{2}/J_{1}=0.5$ (a) and Fermi-Hubbard model on the $8\times 8$ lattice at $U/t=8$ with hole doping $n_{h}=1/8$ (b). Reference energies $E_{\rm ref}$ taken from PEPS $D=10$ ($J_{1}-J_{2}$, $E=-0.490478(4)$) and DMRG ($D=10000$ with SU(2) symmetry, Hubbard model, $E=-0.694391$). The time cost is for computing $\langle\vec{n}|\Psi(D,\chi)\rangle$ for a given $|\vec{n}\rangle$ on a single core. Highlighted symbols (connected by a black line) indicate the optimal $(D,\chi)$ pairs that have smallest time cost for a given accuracy. Here $D=4-8$ and $\chi=2,3,4,5,6,8,10,12$ is used for the $J_{1}-J_{2}$ model and $D=6-12$ and $\chi=2,3,4,5,6,8,10,12,14$ for the Hubbard model.
Refer to captionRefer to caption
@@
As discussed TNFs only require a consistent and efficient contraction of $\Psi(\vec{n})$, and there exist TNFs for which this is the case without any (amplitude dependent) isometries, but for which the efficient contraction of $\langle\Psi|\hat{O}|\Psi\rangle$ is still not possible. (We note that tensor networks where computing $\langle\Psi|\hat{O}|\Psi\rangle$ is efficient but $\Psi(\vec{n})$ is not are used in the form of the multi-scale entanglement renormalization ansatz [51]). Here we describe one such application of TNFs without amplitude dependent isometries.
It has previously been observed that common tensor networks that can be efficiently exactly contracted for $\Psi(\vec{n})$ and $\langle\Psi|\hat{O}\Psi\rangle$, i.e. MPS and tree tensor network states, can be mapped onto efficient neural networks by constructing polynomial size neural nets that emulate tensor network contraction, establishing that exactly contractible tensor networks are a formal subclass of efficient neural quantum states [36]. It has also been suggested that neural networks may not have an efficient tensor network representation, because the well known contractible tensor networks satisfy area laws, while neural networks need not do so [36].
However, if one allows for general tensor network geometries, it is straightforward to construct a TNF (without using isometries) that represents a neural network quantum state without the restriction of an area law. We consider the explicit example of a feed forward neural network (FNN), a popular universal architecture with $k$-layers of neurons, where the outputs of each layer $\{y^{(k-1)}_{i}\}$ form the inputs $\{x^{(k)}_{i}\}$ to the next layer, and the $i$-th neuron (activation) function in layer $k$ is $f^{(k)}_{i}(\{x^{(k)}_{j}\};\theta_{i}^{(k)})$ where $\theta_{i}^{(k)}$ are the neuron parameters. Assuming that all $f$ are implemented without recursion, then the computational graph of a FNN is a directed acyclic graph. Then we observe that (i) if the FNN is efficient, the classical binary circuit implementation of the neural network function is of polynomial size in logic and copy gates, where a binary logic gate is a tensor with two binary input legs and one binary output leg, while copy takes one binary input and returns two binary outputs (nonlinearity of the activation functions enters through the copy operator) (ii) given the input product state $|\vec{n}\rangle$ (i.e. a bitstring), the application of each gate $G$ produces another product state $|\vec{n}^{\prime}\rangle$ (i.e. another bitstring). Thus contracting the resulting tensor network in the appropriate order from inputs to the outputs (for example, using sparse tensor contraction, or by introducing SVDs to discover the product state structure of $|\vec{n^{\prime}}\rangle$, see SM for more details [46]) can always be done efficiently.

### v3

@@ Introduction.
Tensor networks represent the amplitude of a quantum state on an $L$ site lattice, denoted $\langle{\bf n}|\Psi\rangle\equiv\Psi({\bf n})\equiv\Psi(n_{1},n_{2},\ldots,n_{L})$, as the contraction of a set of tensors. As a simple example, we consider $L$ tensors connected in a graph (Fig. 1a), each associated with a lattice site Hilbert space; this is known as a projected entangled pair state (PEPS) [27]. Each tensor contains $dD^{m}$ elements ($m$ is the number of bonds which connect tensors to adjacent tensors): $d$ is the local Hilbert space dimension and $D$ controls the expressivity. For tree graphs, the resulting contraction over the shared tensor bonds can be computed exactly in $O(L)\mathrm{poly}(D)$ time [28]. For general graphs, however, both $\langle{\bf n}|\Psi\rangle$ and $\langle\Psi|\hat{O}|\Psi\rangle$ cannot be exactly contracted efficiently: the cost is exponential in the tree-width of the graph [21]. Approximate contraction algorithms have thus been introduced whereby during the contraction process, the large intermediate tensors that are formed are compressed, via singular value decomposition (SVD), to a controlled size, $O(\mathrm{poly}(\chi))$, where $\chi$ is termed the auxiliary bond dimension [2, 29]. Only in the limit of $\chi\to\infty$ is the contraction exact. The resulting approximate contraction has a cost $O(L)\mathrm{poly}(D)\mathrm{poly}(\chi)$, with an error with respect to the exact contraction controlled by $\chi$.
@@
From the perspective of approximate contraction, tensor network computations should seek to be converged with respect to $\chi$, otherwise the results are not meaningful. For example, for a Hamiltonian $\hat{H}$, if $\langle\Psi|\hat{H}|\Psi\rangle/\langle\Psi|\Psi\rangle$ is computed by approximate contraction, with insufficient $\chi$ the energy may violate the variational theorem. The requirement that computations should use a sufficiently large $\chi$ (such that the evaluation of $\langle\Psi|\hat{O}|\Psi\rangle$ or $\langle{\bf n}|\Psi\rangle$ is no longer changing significantly with $\chi$) places restrictions on the types of graphs that can be used to define the tensor network. For example, in many-body physics simulations, it is believed that only highly structured graphs with limited entanglement, such as (near-)area law states, support an efficient contraction scheme. Because of the latter requirement, it has been argued that, in practical use, tensor networks cannot replicate the expressivity of other flexible types of many-body ansatz, such as neural network quantum states [30, 31, 32, 33, 34, 35, 36].
@@
In the current work, we observe that the above limitations on tensor network computations arise due to an assumption, namely that the tensor network ansatz is only meaningfully defined in the exact contraction limit. We can, however, adopt an alternative perspective, namely, to consider the approximate contraction scheme for any $\chi$ as itself part of the tensor network ansatz for the amplitude $\Psi({\bf n})$. This view is natural and in some ways obvious when using tensor networks in variational Monte Carlo (VMC) calculations [37, 38, 39, 40, 41], but we will argue that this is not a manifestation of tensor networks in a particular numerical algorithm, but rather a change of perspective that alters and broadens the definition and utility of tensor networks. We refer to the larger set of parameterized many-body states that arise from “consistent” contraction schemes (to be defined later) as tensor network functions. This wider class of states removes many restrictions that are assumed to apply to tensor networks. We demonstrate that tensor network functions provide strictly variational estimates of the energy, allow for practical tensor network computations on graphs that are highly challenging to contract accurately, including those that describe volume law physics, and expand the set of states that can be efficiently represented by tensor networks to cover other classes of ansatz, such as neural network quantum states.
@@
Figure 1: (a) PEPS on a general graph with $L$ site tensors. (b) PEPS on a $4\times 4$ square lattice. (c) Insertion of a pair of isometries for approximate tensor contractions; the isometries can be obtained with standard tensor network techniques by SVD [2, 3]. (d) Approximate PEPS contractions by inserting isometries for a given configuration $|{\bf n}\rangle$ (gray circles). (e) For a given input vector $|{\bf n}\rangle=(n_{1},n_{2},...,n_{L})$, a tensor network function outputs a unique value $\Psi({\bf n})$.
@@
A tensor network function (TNF) is a consistent contraction of the tensor network representation of the amplitude $\Psi({\bf n})$. To illustrate the basic idea, we first consider a standard PEPS for a two-dimensional lattice, illustrated in Fig. 1b. Because the graph has loops, the PEPS amplitude cannot be computed efficiently according to Fig. 1b. Instead, in practice, the amplitude is defined by a different tensor network illustrated in Fig. 1d. In this figure, isometries (triangles) with bond dimension $\chi$ have been inserted, where the values of the isometries are determined by an SVD of the contraction of neighbouring tensors (see Fig. 1c) [2, 3]. The final amplitude is obtained by contracting the tensors and isometries together. The pattern of isometries used here corresponds to a specific approximate contraction method (boundary contraction [27, 2]) and other choices can be made [29, 42], but the essential purpose is to make the contraction of the modified graph containing the original tensors and the isometries efficient.
Note, however, that although the structure of the tensor network with isometries in Fig. 1d now appears pseudo-one-dimensional, this does not mean that its entanglement structure (for $\chi>1$) is reduced to that of a matrix product state (MPS). This is because the isometries are themselves functions of the configuration $|{\bf n}\rangle$, i.e. different values of the isometries are obtained, via SVD, for different configurations. However, we choose to keep the same positions of the isometries in the graph for any configuration (even though the entries are different); we refer to this as using fixed (position) isometries, which defines a consistent contraction. Because of the configuration dependence of the isometry entries, the graph with isometries no longer represents a multi-linear ansatz like the original tensor network state; it is only a multi-linear ansatz in the limit that $\chi\to\infty$ (and the isometries become identities). Instead, it defines a tensor network function: for any given $(D,\chi)$, any $|{\bf n}\rangle$ is mapped to a single value, the amplitude, through an efficient contraction. The efficient TNF representation of $\Psi({\bf n})$ does not guarantee deterministic efficiency of computing $\langle\Psi|\hat{O}|\Psi\rangle$; but such expectations may be evaluated stochastically by sampling with efficient sample complexity.
We emphasize that here we are defining a new class of state i.e. the TNF, whose amplitudes are exactly given by a certain contraction graph of a tensor network. The essential differences between TNF and conventional tensor network computations are further illustrated in the following examples, and discussed also in Section-1 in SM [43].
@@
VMC algorithms also compute $\Psi({\bf n})$ via approximate contraction (for a brief review of tensor network VMC (TN-VMC) algorithms, see SM [43] or Refs. [37, 38, 39, 40, 41]) and it is instructive to consider the differences between the standard TN-VMC approximation of $\Psi({\bf n})$ and the definition of a TNF above. There are two seemingly small, but significant, differences (i) in TN-VMC, $\chi$ is viewed as a parameter (typically $\chi=\mathrm{const}\times D$, with $\mathrm{const}>1$) where $\chi$ is increased to convergence for a fixed $D$, and (ii) the positions of the isometries are not fixed but dependent on the amplitude that is being computed (we refer to this as dynamic isometries), in order to allow for the reuse of computations when computing amplitudes of configurations which differ by only a few substitutions [41]. See SM for details [43].
The most important distinction from TNF is that the use of dynamic isometries in TN-VMC does not lead to a consistent function. In other words, given the same configuration $|{\bf n}\rangle$, $\Psi({\bf n})$ needs not be the same, as it depends on the prior configuration and which components of the tensor network contraction are being reused. Only in the limit of $\chi\to\infty$ are the amplitudes consistent and the function defined, which underlies the usual need to converge $\chi$ in TN-VMC calculations. In contrast, it is clear that with using fixed isometries, the corresponding TNF is consistent for any $D$ and any $\chi$. See SM for details [43].
@@
In Fig. 2(a) we show the energy of an optimized PEPS ($D=8$, various $\chi$) TNF for the ground-state of the open-boundary condition (OBC) spin-1/2 square-lattice $32\times 32$ Heisenberg model defined using the boundary contraction computational graph, see SM, compared to that of a PEPS in standard TN-VMC using dynamic isometries, where $\chi$ controls the error of approximate contraction [41]. We see immediately that the PEPS TNF energy is variational for all $\chi$. In contrast, viewing the PEPS amplitude as an (inconsistent) approximate contraction as in standard VMC leads to a non-variational energy, where the non-variationality is typically regarded as the “contraction error”. This behaviour is even more clear for a lattice that is usually considered more difficult to contract. Using the square-lattice $J_{1}-J_{2}$ model with periodic boundary conditions (PBC) as an example, in Fig. 2(b) we show the same comparison between TNF and standard TN-VMC. In addition to the variational versus non-variational nature of the energies, we also see a large variance of the TN-VMC energy arising from the inconsistent definition of amplitudes. This leads to substantial difficulties in the stochastic optimization of the ansatz. Overall, we see that lack of variationality is not intrinsic to approximate contraction: it results from the perspective on the definition of the ansatz.
@@
Figure 3: Time cost (in seconds) versus accuracy of the tensor network function $|\Psi(D,\chi)\rangle$, for the $J_{1}-J_{2}$ model on the $16\times 16$ lattice at $J_{2}/J_{1}=0.5$ (a) and Fermi-Hubbard model on the $8\times 8$ lattice at $U/t=8$ with hole doping $n_{h}=1/8$ (b). Reference energies $E_{\rm ref}$ taken from PEPS $D=10$ ($J_{1}-J_{2}$, $E=-0.490478(4)$) and DMRG ($D=10000$ with SU(2) symmetry, Hubbard model, $E=-0.694391$). The time cost is for computing $\langle{\bf n}|\Psi(D,\chi)\rangle$ for a given $|{\bf n}\rangle$ on a single core. Highlighted symbols (connected by a black line) indicate the optimal $(D,\chi)$ pairs that have smallest time cost for a given accuracy. Here $D=4-8$ and $\chi=2\sim 12$ is used for the $J_{1}-J_{2}$ model and $D=6-12$ and $\chi=2\sim 14$ for the Hubbard model. The insets present the energy variation of the TNF $|\Psi(D,\chi)\rangle$ with $D$ and $\chi$.
@@
In Figs. 3(a) and (b), we show results from approximating the ground-state of a spin-$1/2$ square-lattice $16\times 16$ $J_{1}-J_{2}$ Heisenberg model for $J_{2}/J_{1}=0.5$, and an $8\times 8$ Hubbard model with $U/t=8$ and $1/8$ hole doping, using bosonic and fermionic PEPS TNFs for OBC systems [27, 44, 45, 46], respectively. We minimize the energy for many different $(D,\chi)$ combinations using gradient descent, and we plot the computational cost of the amplitude against the obtained accuracy. For a specified relative accuracy, we can then search for the $(D,\chi)$ pair with the lowest computational cost. We see that contrary to the standard situation with tensor network states where one uses $\chi>D$, the most powerful TNFs for a given computational cost have $\chi<D$. For example, for $\sim 0.01$ relative accuracy in the Hubbard model, using $(D,\chi)=(12,6)$ yields the same expressivity as $(D,\chi)=(10,10)$, but with lower cost. In no case is it computationally more efficient to use $\chi>D$. The insets of Fig. 3 presenting the minimized energy for every ansatz $|\Psi(D,\chi)\rangle$, show that the most rapid improvement in energy is obtained at small $\chi$, rather than saturating $\chi$ for a given $D$, confirming the above conclusion.
@@
As discussed TNFs only require a consistent and efficient contraction of $\Psi({\bf n})$, and there exist TNFs for which this is the case without any (amplitude dependent) isometries, but for which the efficient contraction of $\langle\Psi|\hat{O}|\Psi\rangle$ is still not possible. (We note that tensor networks where computing $\langle\Psi|\hat{O}|\Psi\rangle$ is efficient but $\Psi({\bf n})$ is not are used in the form of the multi-scale entanglement renormalization ansatz [51]). Here we describe one such application of TNFs without amplitude dependent isometries.
It has previously been observed that common tensor networks that can be efficiently exactly contracted for $\Psi({\bf n})$ and $\langle\Psi|\hat{O}|\Psi\rangle$, i.e. MPS and tree tensor network states, can be mapped onto efficient neural networks by constructing polynomial size neural nets that emulate tensor network contraction, establishing that exactly contractible tensor networks are a formal subclass of efficient neural quantum states [36]. It has also been suggested that neural networks may not have an efficient tensor network representation, because the well known contractible tensor networks satisfy area laws, while neural networks need not do so [36].
However, if one allows for general tensor network geometries, it is straightforward to construct a TNF (without using isometries) that represents a neural network quantum state without the restriction of an area law. We consider the explicit example of a feed forward neural network (FNN), a popular universal architecture with $k$-layers of neurons, where the outputs of each layer $\{y^{(k-1)}_{i}\}$ form the inputs $\{x^{(k)}_{i}\}$ to the next layer, and the $i$-th neuron (activation) function in layer $k$ is $f^{(k)}_{i}(\{x^{(k)}_{j}\};\theta_{i}^{(k)})$ where $\theta_{i}^{(k)}$ are the neuron parameters. Assuming that all $f$ are implemented without recursion, then the computational graph of a FNN is a directed acyclic graph. Then we observe that (i) if the FNN is efficient, the classical binary circuit implementation of the neural network function is of polynomial size in logic and copy gates, where a binary logic gate is a tensor with two binary input legs and one binary output leg, while copy takes one binary input and returns two binary outputs (nonlinearity of the activation functions enters through the copy operator) (ii) given the input product state $|{\bf n}\rangle$ (i.e. a bitstring), the application of each gate $G$ produces another product state $|{\bf n}^{\prime}\rangle$ (i.e. another bitstring). Thus contracting the resulting tensor network in the appropriate order from inputs to the outputs (for example, using sparse tensor contraction, or by introducing SVDs to discover the product state structure of $|{\bf n}^{\prime}\rangle$, see SM for more details [43]) can always be done efficiently.